Ali is trying to find the limit of a function which is expressed as L = lim x→0+ (1 + x)1/x . From his understanding, the quantity (1 + x), must be greater than 1 for x > 0. Furthermore, the power1 / x is going to infinity as ? approaches 0 from the right. So L is the result of taking a number greater than 1 to higher power, therefore L = ∞. On the other hand, he sees that (1 + x) is approaching 1 as x approaches 0, and 1 taken to any power whatever is 1. Ali concluded, L = 1. Help Ali by pointing out to him the error of his ways.


1
Expert's answer
2020-08-31T17:15:34-0400

Let "x_n=\dfrac{1}{n}, n\in\Z^+"


"\lim\limits_{x_n\to 0^+}(1+x_n)^{{1 \over x_n}}=\lim\limits_{n\to \infin}(1+{1\over n})^{n}"

Use Newton's Binomial formula

"(1+{1\over n})^{n}=1+n\cdot{1\over n}+{n(n-1)\over 1\cdot2}\cdot{1\over n^2}+"

"+{n(n-1)(n-2)\over 1\cdot2\cdot3}\cdot{1\over n^3}+...+{n(n-1)...(1)\over 1\cdot2\cdot...\cdot n}\cdot{1\over n^n}>1+1=2"



Let "f(x)=\ln (x)," then "f'(x)=1\/x." Thus "f'(1)=1\/1=1."

From the definition of a derivative as a limit, we have


"f'(1)=\lim\limits_{h\to 0}\dfrac{f(1+h)-f(1)}{h}="

"=\lim\limits_{x\to 0}\dfrac{f(1+x)-f(1)}{x}=\lim\limits_{x\to 0}\dfrac{\ln(1+x)-
\ln(1)}{x}="

"=\lim\limits_{x\to 0}\dfrac{\ln(1+x)
}{x}=\lim\limits_{x\to 0}\ln(1+x)^{{1\over x}}"

Because "f'(1)=1," we have


"\lim\limits_{x\to 0}\ln(1+x)^{{1\over x}}=1"

Theorem

If "f" is continuous at "b" and "\lim\limits_{x\to a}g(x)=b," then "\lim\limits_{x\to a}f(g(x))=f(b)"


"\lim\limits_{x\to a}f(g(x))=f(\lim\limits_{x\to a}g(x))"

Since the exponential function is continuous, we have


"e=e^1=e^{\lim\limits_{x\to 0}\ln(1+x)^{{1\over x}}}=\lim\limits_{x\to 0}e^{\ln(1+x)^{{1\over x}}}="

"=\lim\limits_{x\to 0}(1+x)^{{1\over x}}"

Therefore


"\lim\limits_{x\to 0}(1+x)^{{1\over x}}=e"

"\lim\limits_{x\to 0^+}(1+x)^{{1\over x}}=e"